OpenStudy (anonymous):

find an equation for the line passing through the point (1,2) and parallel to the line y=3x-2

OpenStudy (anonymous):

what do you know about parallel lines and their slopes

OpenStudy (anonymous):

not much

OpenStudy (anonymous):

alright you know that slope = rise/run correct or it's the change of y over change of x now if both are parallel 

OpenStudy (campbell_st):

the gradient of the given line is m = 3 Parallel lines have the same slopes or gradients. so the line is in the form y = 3x + b just substitute the values from the point x = 1 and y = 2 to find the value of b, the y-intercept. then you'll have your line equation.

OpenStudy (anonymous):

In this picture you can see both lines are parallel. if you find the slope of each line you'll find that as x increases and decreases y stays the same this means the change of y =0 and change of x is anything you want it to be. in this case x=1 so m=0/1 for both lines. Parallel lines always have the same slope

OpenStudy (anonymous):

knowing this if you want to make a line parallel to another line y=mx+b, you'll know that both lines will have slope m. then all you need to do is use the point slope form of a line \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

where \[x_1,y_1\] is a point on the line and m is the slope

OpenStudy (anonymous):

so is my answer y=3x-1?

OpenStudy (anonymous):

what do i do if its perpendicular?